# Robust covariance estimation and explainable outlier detection for matrix-valued data

### v1

@@ 7.3 Video data
For our analyis, we converted the original RGB video to a grayscale video, applied the Fast-MMCD procedure, and obtained MMDs for all 633 frames, which are visualized in Figure 8. The plot on the left shows the robust MMDs for all 633 frames, and the one on the right for frames 471 to 633 to better highlight the increase in MMD when the man enters the scenery, with a short drop in MMD when he disappears behind the tree. We indicate frames 487, 491, and 495, also presented in Figure 9 in terms of their cellwise Shapley values. We see that the pixels that form the contours of the man and most of the pixels of the man’s head contribute most to the outlyingness. When the man disappears behind the tree, there are fewer pixels with high outlyingness contributions. Since the sum of the contributions amounts to the squared MMD of an observation, this explains the behavior of the MMDs of the frames shown in Figure 7(b).
@@ 8 Summary and conclusions
This work introduced the MMCD (matrix minimum covariance determinant) estimators as a robust counterpart to the ML estimators in the matrix-variate normal model. Several desirable properties are achieved: equivariance under linear matrix transformations, high breakdown point, and consistency under elliptical matrix-variate distributions. The proposed reweighted versions lead to higher efficiency but not to any loss in terms of breakdown point. An algorithm along the lines of the FastMCD procedure (Rousseeuw and Driessen,, 1999) allows for efficient computation of the estimators. Simulation experiments validate the theoretical properties and advantages over robust estimation for vectorized observations regarding breakdown and computational efficiency.
@@
The R package robustmatrix available on CRAN includes a parallelized C++ implementation of the FastMMCD algorithm and a vignette to reproduce the examples presented in this paper.
Supplementary material to:
Robust covariance estimation and explainable outlier detection for matrix-valued data
Marcus Mayrhofer, Una Radojicic, and Peter Filzmoser
@@ Appendix A Preliminaries
respectively. The computation of the MLEs for matrix-variate samples based on Equations (2)-(4) involves estimating $p(p+1)/2+q(q+1)/2$ parameters instead of $pq(pq+1)/2+1$ parameters for the vectorized observations according to Equation (33). This raises the question of whether fewer than $pq+1$ observations are sufficient for guaranteeing the existence and uniqueness of MLEs for i.i.d. samples from a matrix normal distribution. This question was investigated in several papers, such as Dutilleul, (1999); Lu and Zimmerman, (2005); Srivastava et al., (2008); Roś et al., (2016); Soloveychik and Trushin, (2016). We rely on the latter for the most recent proof of those conditions. Note that it is not necessary to assume that the sample consists of i.i.d. observations. In fact, the i.i.d. assumption can be relaxed to allow for statistically dependent samples and it is not even necessary to require identical distribution (Soloveychik and Trushin,, 2016, Remarks 2 and 6). The critical condition for existence and uniqueness is that the sample contains at least $n\geq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$ observations that are not collinear. The same holds for the existence and uniqueness of the MMCD estimators, where $n$ is replaced by $h$, and for properties like the breakdown point the assumptions could be relaxed only requiring that the sample is in general position, i.e., no subset of $r,2\leq r\leq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$ samples lies on an $r-2$ dimensional subspace. However, the i.i.d. assumption is still necessary when we consider properties like consistency.
The idea of the multivariate MCD estimator is as follows: Let $\bm{x}_{i}=(x_{i1},\ldots,x_{ip})^{\prime}\in\mathbb{R}^{p}$ denote the $i$-th observation of a data set in the multivariate setting, where $i=1,\ldots,n$. The objective of the MCD estimator is to find the subset of $h$ out of $n$ observations whose sample covariance matrix has the lowest determinant, with $\nicefrac{{n}}{{2}}\leq h\leq n$ and $h>p$. In total, there are $\binom{h}{n}$ possible $h$-subsets, and thus, a strategy needs to be used to tackle the optimization problem efficiently. This has been done with the so-called Fast-MCD algorithm (Rousseeuw and Driessen,, 1999), which internally sorts the observations based on their Mahalanobis distances. For an observation $\bm{x}_{i}$ from a population with mean $\bm{\mu}\in\mathbb{R}^{p}$ and covariance $\bm{\Sigma}\in\operatorname{PSD}(p)$ it is given by
@@ B.1 Properties of MMCD estimators
be the ellipsoid containing the matrices $\bm{X}\in\mathbb{R}^{p\times q}$ with $\operatorname{MMD}^{2}(\bm{X};\bm{T},\bm{U},\bm{V})\leq 1$, where $\bm{T}\in\mathbb{R}^{p\times q}$, $\bm{U}\in\operatorname{PSD}(p)$ and $\bm{V}\in\operatorname{PSD}(q)$. The volume of this ellipsoid is given by
@@
If $\bm{A}\in\operatorname{PSD}(p)$ with eigenvalues $0<\lambda_{p}(\bm{A})\leq\ldots\leq\lambda_{1}(\bm{A})$ then the eigenvalues of $\bm{A}^{-1}$ are the reciprocals of the eigenvalues of $\bm{A}$, i.e. $\lambda_{i}(\bm{A}^{-1})=\lambda^{-1}_{i}(\bm{A})$. Hence, we have that
@@
Suppose $\bm{A}\in\operatorname{PSD}(p)$, $\bm{B}\in\operatorname{PSD}(q)$ and let $\lambda(\bm{A})$ be an eigenvalue of $\bm{A}$ with corresponding eigenvector $\bm{v}(\bm{A})$, and $\lambda(\bm{B})$ an eigenvalue of $\bm{B}$ with corresponding eigenvector $\bm{v}(\bm{B})$. Then $\lambda(\bm{A})\lambda(\bm{B})$ is an eigenvalue of $\bm{B}\otimes\bm{A}$ with corresponding eigenvector $\bm{v}(\bm{B})\otimes\bm{v}(\bm{A})$. We denote the sequence of eigenvalues of $\bm{A}$ and $\bm{B}$ as $0<\lambda_{p}(\bm{A})\leq\ldots\leq\lambda_{1}(\bm{A})$ and $0<\lambda_{q}(\bm{b})\leq\ldots\leq\lambda_{1}(\bm{b})$, respectively. It follows that the smallest eigenvalue $\lambda_{pq}(\bm{A},\bm{B})=\lambda_{p}(\bm{A})\lambda_{q}(\bm{B})$ and the largest eigenvalue $\lambda_{1}(\bm{A},\bm{B})=\lambda_{1}(\bm{A})\lambda_{1}(\bm{B})$. Moreover note that for $\bm{Z}\in\mathbb{R}^{p\times q}$
@@
subject to $w_{1},\dots,w_{n}\in\{0,1\}$, $\sum_{i=1}^{n}w_{i}=h$, $\hat{\bm{\mu}}\in\mathbb{R}^{pq}$, $\hat{\bm{\Sigma}}\in\operatorname{PSD}(pq)$; see Raymaekers and Rousseeuw, (2023) for more insight. Similarly, the MMCD estimator is a solution to the following maximization problem:
@@
subject to $w_{1},\dots,w_{n}\in\{0,1\}$, $\sum_{i=1}^{n}w_{i}=h$, $\hat{\bm{\operatorname{M}}}\in\mathbb{R}^{p\times q}$, $\hat{\bm{\Sigma}}^{\operatorname{row}}\in\operatorname{PSD}(p)$, $\hat{\bm{\Sigma}}^{\operatorname{col}}\in\operatorname{PSD}(q)$; see Proposition 2.0.1.
Denote further $(\bm{w}_{\operatorname{MCD}},\hat{\bm{\mu}}_{\operatorname{MCD}},\hat{\bm{\Sigma}}_{\operatorname{MCD}})$ and $(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}})$ weights, mean and covariance estimators for the vectorized sample $(\bm{x}_{1},\dots,\bm{x}_{n})$, based on MCD and MMCD, respectively. As $\bm{X}_{i}\sim\mathcal{MN}(\bm{\operatorname{M}},\bm{\Sigma},\bm{\Omega})$, then $\bm{x}_{i}\sim\mathcal{N}(\bm{\mu},\bm{\Omega}\otimes\bm{\Sigma})$, with $\bm{\mu}=\operatorname{vec}(\bm{\operatorname{M}})$ and $i=1,\dots,n$. Moreover, the mean estimator $\hat{\bm{\mu}}_{\operatorname{MCD}}$ and properly scaled covariance estimator $\hat{\bm{\Sigma}}_{\operatorname{MCD}}$ are strongly consistent for the population counterparts $\bm{\mu}$ and $\bm{\Omega}\otimes\bm{\Sigma}$; see e.g. Croux and Haesbroeck, (1999) and Cator and Lopuhaä, (2012). Especially, this implies that for every $\delta>0$ there exists $n\in\mathbb{N}$ such that
@@
for some $\bm{A}\otimes\bm{B}\in\operatorname{PSD}(p)\otimes\operatorname{PSD}(q)$. In the following, we will drop ${a.s.}$ superscript from (in)equality signs when it is clear from the context. For fixed weights $\bm{w}$, the log-likelihood function $l_{\cdot,\bm{w}|(\bm{x}_{1},\dots,\bm{x}_{n})}:(\bm{\mu},\bm{\Sigma})\mapsto l(\bm{\mu},\bm{\Sigma}|\bm{w},(\bm{x}_{1},\dots,\bm{x}_{n}))$ is continuous in both $\bm{\mu}$, and $\bm{\Sigma}$, and its continuity implies
@@
Similarly, $(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}})$ is a maximizer of $l(\cdot|(\bm{x}_{1},\dots,\bm{x}_{n}))$ in the set of all feasible weights, means, and covariances with Kronecker product structure. As $(\bm{w}_{\operatorname{MCD}},\bm{\mu},\bm{A}\otimes\bm{B})$ belongs to the same set,
$l(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}|(\bm{x}_{1},\dots,\bm{x}_{n}))>l(\bm{w}_{\operatorname{MCD}},\bm{\mu},\bm{A}\otimes\bm{B}|(\bm{x}_{1},\dots,\bm{x}_{n})).$
Denote further $\hat{\bm{S}}_{\operatorname{MMCD}}=\frac{1}{h}\sum_{i=1}^{n}w_{{\operatorname{MMCD}},i}(\bm{x}_{i}-\hat{\bm{\mu}}_{\mathrm{MMCD}})(\bm{x}_{i}-\hat{\bm{\mu}}_{\mathrm{MMCD}})^{\prime}$ to be the estimate of the covariance of the vectorized observations, based on the weights (subset) produced by the MMCD algorithm. As $\hat{\bm{S}}_{\operatorname{MMCD}}$ is optimal for $l$ given fixed weights $\bm{w}_{\mathrm{MMCD}}$,
@@
$>l(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}|(\bm{x}_{1},\dots,\bm{x}_{n}))$
@@
$\hat{\bm{\mu}}_{\operatorname{MMCD}}\xrightarrow{a.s.}\bm{\mu},\quad\hat{\bm{S}}_{\operatorname{MMCD}}\xrightarrow{a.s.}c_{1}^{-1}\bm{\Omega}\otimes\bm{\Sigma}\quad,$ (46)
where $c_{1}>0$ is a distribution-specific consistency factor of MCD for the given in Croux and Haesbroeck, (1999).
To complete the proof consider reparametrization of $l(\bm{w},\bm{a},\bm{A}|(\bm{x}_{1},\dots,\bm{x}_{n}))$ for fixed weights $\bm{w}\in\mathbb{R}^{n}$, mean $\bm{a}\in\mathbb{R}^{pq}$, and covariance $\bm{A}\in\operatorname{PSD}(pq)$ in terms of the precicion matrix $\bm{B}=\bm{A}^{-1}$. Denote this new parametrization as $g(\bm{B}|\bm{w},\bm{a},(\bm{x}_{1},\dots,\bm{x}_{n}))=l(\bm{w},\bm{a},\bm{B}^{-1}|\bm{x}_{1},\dots,\bm{x}_{n})$, which is now concave in $\bm{B}$. Especially, for $\bm{w}=\bm{w}_{\operatorname{MMCD}}$ and $\bm{a}=\hat{\bm{\mu}}_{\operatorname{MMCD}}$, the function $g(\bm{B}|\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},(\bm{x}_{1},\dots,\bm{x}_{n}))$ is concave in $\bm{B}$ and achieves a unique global maximum at $\bm{B}=\hat{\bm{S}}_{\operatorname{MMCD}}$. Equations (43) and (B.1) then give
@@
$-l(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}|(\bm{x}_{1},\dots,\bm{x}_{n}))<\varepsilon,$
@@
$-g((\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}})^{-1}|\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},(\bm{x}_{1},\dots,\bm{x}_{n}))<\varepsilon,$
as both $\hat{\bm{S}}_{\operatorname{MMCD}}$ and $\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}$ are a.s. positive definite for $n$ large enough. Concavity of $g$ and the fact that $\hat{\bm{S}}_{\operatorname{MMCD}}^{-1}$ is its global maximum further imply that
$\|\hat{\bm{S}}_{\operatorname{MMCD}}^{-1}-(\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}})^{-1}\|<\delta_{1},$
for $\delta_{1}=\delta_{1}(\varepsilon)\to 0$ as $n\to\infty$. Almost sure positive definiteness of $\hat{\bm{S}}_{\operatorname{MMCD}}$ and $\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}$, and continuity of matrix inverse imply that
$\|\hat{\bm{S}}_{\operatorname{MMCD}}-\hat{\bm{\Omega}}_{\operatorname{MMCD}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}\|<\delta,$ (47)

### v2

@@ 7.3 Video data
For our analyis, we converted the original RGB video to a grayscale video, applied the MMCD procedure, and obtained MMDs for all 633 frames, which are visualized in Figure 9. The plot on the left shows the robust MMDs for all 633 frames, and the one on the right for frames 471 to 633 to better highlight the increase in MMD when the man enters the scenery, with a short drop in MMD when he disappears behind the tree. We indicate frames 487, 491, and 495, also presented in Figure 10 in terms of their cellwise Shapley values. We see that the pixels that form the contours of the man and most of the pixels of the man’s head contribute most to the outlyingness. When the man disappears behind the tree, there are fewer pixels with high outlyingness contributions. Since the sum of the contributions amounts to the squared MMD of an observation, this explains the behavior of the MMDs of the frames shown in Figure 9(b).
@@ 8 Summary and conclusions
This work introduced the MMCD (matrix minimum covariance determinant) estimators as a robust counterpart to the ML estimators in the matrix-variate normal model. Several desirable properties are achieved: equivariance under matrix affine transformations, high breakdown point, and consistency under elliptical matrix-variate distributions. The proposed reweighted versions lead to higher efficiency but not to any loss in terms of breakdown point. An algorithm along the lines of the Fast-MCD procedure (Rousseeuw and Driessen,, 1999) allows for efficient computation of the estimators. Simulation experiments validate the theoretical properties and advantages. Depending on the ratio of the number of rows and columns of the matrix-valued observations, the MMCD estimators show a big advantage over robust estimation for vectorized observations regarding breakdown and computational efficiency.
@@
The R package robustmatrix includes a parallelized C++ implementation of the MMCD algorithm and a vignette to reproduce the examples presented in this paper.
Acknowledgements
This work was supported by the AI4CSM project and has received funding from the ECSEL Joint Undertaking (JU) under grant agreement No 101007326; and the Austrian IKT der Zukunft programme via the Austrian Research Promotion Agency (FFG) and the Austrian Federal Ministry for Climate Action, Environment, Energy, Mobility, Innovation and Technology (BMK) under project No 884070. This work was supported by the Austrian Science Fund (FWF), project number I 5799-N.
@@ Appendix A Preliminaries
respectively. The computation of the MLEs for matrix-variate samples based on Equations (3)-(5) involves estimating $p(p+1)/2+q(q+1)/2+pq$ parameters instead of $pq(pq+1)/2+pq$ parameters for the vectorized observations according to Equation (32). This raises the question of whether fewer than $pq+1$ observations are sufficient for guaranteeing the existence and uniqueness of MLEs for i.i.d. samples from a matrix normal distribution. This question was investigated in several papers, such as Dutilleul, (1999); Lu and Zimmerman, (2005); Srivastava et al., (2008); Roś et al., (2016); Soloveychik and Trushin, (2016). We rely on the latter for the most recent proof of those conditions. Note that it is not necessary to assume that the sample consists of i.i.d. observations. In fact, the i.i.d. assumption can be relaxed to allow for statistically dependent samples and it is not even necessary to require identical distribution (Soloveychik and Trushin,, 2016, Remarks 2 and 6). The critical condition for existence and uniqueness is that the sample contains at least $n\geq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$ observations that are not collinear. The same holds for the existence and uniqueness of the MMCD estimators, where $n$ is replaced by $h$, and for properties like the breakdown point the assumptions could be relaxed only requiring that the sample is in general position, i.e., no subset of $r,2\leq r\leq\lfloor\nicefrac{{p}}{{q}}+\nicefrac{{q}}{{p}}\rfloor+2$ samples lies on an $r-2$ dimensional subspace. However, the i.i.d. assumption is still necessary when we consider properties like consistency.
The idea of the multivariate MCD estimator is as follows: Let $\bm{x}_{i}=(x_{i1},\ldots,x_{ip})^{\prime}\in\mathbb{R}^{p}$ denote the $i$-th observation of a data set in the multivariate setting, where $i=1,\ldots,n$. The objective of the MCD estimator is to find the subset of $h$ out of $n$ observations whose sample covariance matrix has the lowest determinant, with $\nicefrac{{n}}{{2}}\leq h\leq n$ and $h>p$. In total, there are $\binom{h}{n}$ possible $h$-subsets, and thus, a strategy needs to be used to tackle the optimization problem efficiently. This has been done with the so-called Fast-MCD algorithm (Rousseeuw and Driessen,, 1999), which internally sorts the observations based on their Mahalanobis distances. For an observation $\bm{x}_{i}$ from a population with mean $\bm{\mu}\in\mathbb{R}^{p}$ and covariance $\bm{\Sigma}\in\operatorname{PDS}(p)$ it is given by
@@ B.1 Properties of MMCD estimators
be the ellipsoid containing the matrices $\bm{X}\in\mathbb{R}^{p\times q}$ with $\operatorname{MMD}^{2}(\bm{X};\bm{T},\bm{U},\bm{V})\leq 1$, where $\bm{T}\in\mathbb{R}^{p\times q}$, $\bm{U}\in\operatorname{PDS}(p)$ and $\bm{V}\in\operatorname{PDS}(q)$. The volume of this ellipsoid is given by
@@
If $\bm{A}\in\operatorname{PDS}(p)$ with eigenvalues $0<\lambda_{p}(\bm{A})\leq\ldots\leq\lambda_{1}(\bm{A})$ then the eigenvalues of $\bm{A}^{-1}$ are the reciprocals of the eigenvalues of $\bm{A}$, i.e. $\lambda_{i}(\bm{A}^{-1})=\lambda^{-1}_{i}(\bm{A})$. Hence, we have that
@@
Suppose $\bm{A}\in\operatorname{PDS}(p)$, $\bm{B}\in\operatorname{PDS}(q)$ and let $\lambda(\bm{A})$ be an eigenvalue of $\bm{A}$ with corresponding eigenvector $\bm{v}(\bm{A})$, and $\lambda(\bm{B})$ an eigenvalue of $\bm{B}$ with corresponding eigenvector $\bm{v}(\bm{B})$. Then $\lambda(\bm{A})\lambda(\bm{B})$ is an eigenvalue of $\bm{B}\otimes\bm{A}$ with corresponding eigenvector $\bm{v}(\bm{B})\otimes\bm{v}(\bm{A})$. We denote the sequence of eigenvalues of $\bm{A}$ and $\bm{B}$ as $0<\lambda_{p}(\bm{A})\leq\ldots\leq\lambda_{1}(\bm{A})$ and $0<\lambda_{q}(\bm{b})\leq\ldots\leq\lambda_{1}(\bm{b})$, respectively. It follows that the smallest eigenvalue $\lambda_{pq}(\bm{A},\bm{B})=\lambda_{p}(\bm{A})\lambda_{q}(\bm{B})$ and the largest eigenvalue $\lambda_{1}(\bm{A},\bm{B})=\lambda_{1}(\bm{A})\lambda_{1}(\bm{B})$. Moreover note that for $\bm{Z}\in\mathbb{R}^{p\times q}$
@@
subject to $w_{1},\dots,w_{n}\in\{0,1\}$, $\sum_{i=1}^{n}w_{i}=h$, $\hat{\bm{\mu}}\in\mathbb{R}^{pq}$, $\hat{\bm{\Sigma}}\in\operatorname{PDS}(pq)$; see Raymaekers and Rousseeuw, (2023) for more insight. Similarly, the MMCD estimator is a solution to the following maximization problem:
@@
subject to $w_{1},\dots,w_{n}\in\{0,1\}$, $\sum_{i=1}^{n}w_{i}=h$, $\hat{\bm{\operatorname{M}}}\in\mathbb{R}^{p\times q}$, $\hat{\bm{\Sigma}}^{\operatorname{row}}\in\operatorname{PDS}(p)$, $\hat{\bm{\Sigma}}^{\operatorname{col}}\in\operatorname{PDS}(q)$; see Proposition 2.0.1.
Denote further $(\bm{w}_{\operatorname{MCD}},\hat{\bm{\mu}}_{\operatorname{MCD}},\hat{\bm{\Sigma}}_{\operatorname{MCD}})$ and $(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}})$ weights, mean and covariance estimators for the vectorized sample $(\bm{x}_{1},\dots,\bm{x}_{n})$, based on MCD and MMCD, respectively. As $\bm{X}_{i}\sim\mathcal{ME}(\bm{\operatorname{M}},\bm{\Sigma}^{\operatorname{row}},\bm{\Sigma}^{\operatorname{col}},g)$, then $\bm{x}_{i}\sim\mathcal{E}(\bm{\mu},\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}},g)$, with $\mathbb{E}(\bm{x}_{i})=\bm{\mu}=\operatorname{vec}(\bm{\operatorname{M}})$, $\operatorname{cov}{(\bm{x}_{i})}=c_{g}\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}}$, where $c_{g}$ is a distribution-specific scaling parameter; for more details see Theorem 2.11 in Gupta and Varga, (2012). Moreover, the mean estimator $\hat{\bm{\mu}}_{\operatorname{MCD}}$ and properly scaled covariance estimator $\hat{\bm{\Sigma}}_{\operatorname{MCD}}$ are strongly consistent for the population counterparts $\bm{\mu}$ and $\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}}$; see e.g. Croux and Haesbroeck, (1999) and Cator and Lopuhaä, (2012). Especially, this implies that for every $\delta>0$ there exists $n\in\mathbb{N}$ such that
@@
for some $\bm{A}\otimes\bm{B}\in\operatorname{PDS}(p)\otimes\operatorname{PDS}(q)$. In the following, we will drop ${a.s.}$ superscript from (in)equality signs when it is clear from the context. For fixed weights $\bm{w}$, the log-likelihood function $l_{\cdot,\bm{w}|(\bm{x}_{1},\dots,\bm{x}_{n})}:(\bm{\mu},\bm{\Sigma})\mapsto l(\bm{\mu},\bm{\Sigma}|\bm{w},(\bm{x}_{1},\dots,\bm{x}_{n}))$ is continuous in both $\bm{\mu}$, and $\bm{\Sigma}$, and its continuity implies
@@
Similarly, $(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}})$ is a maximizer of $l(\cdot|(\bm{x}_{1},\dots,\bm{x}_{n}))$ in the set of all feasible weights, means, and covariances with Kronecker product structure. As $(\bm{w}_{\operatorname{MCD}},\bm{\mu},\bm{A}\otimes\bm{B})$ belongs to the same set,
$l(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}}|(\bm{x}_{1},\dots,\bm{x}_{n}))>l(\bm{w}_{\operatorname{MCD}},\bm{\mu},\bm{A}\otimes\bm{B}|(\bm{x}_{1},\dots,\bm{x}_{n})).$
Denote further $\hat{\bm{S}}_{\operatorname{MMCD}}=\frac{1}{h}\sum_{i=1}^{n}w_{{\operatorname{MMCD}},i}(\bm{x}_{i}-\hat{\bm{\mu}}_{\mathrm{MMCD}})(\bm{x}_{i}-\hat{\bm{\mu}}_{\mathrm{MMCD}})^{\prime}$ to be the estimate of ${\bm{\Sigma}}^{\operatorname{col}}\otimes{\bm{\Sigma}}^{\operatorname{row}}$, based on the weights (subset) produced by the MMCD algorithm. As $\hat{\bm{S}}_{\operatorname{MMCD}}$ is optimal for $l$ given fixed weights $\bm{w}_{\mathrm{MMCD}}$,
@@
$>l(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}}|(\bm{x}_{1},\dots,\bm{x}_{n}))$
@@
$\hat{\bm{\mu}}_{\operatorname{MMCD}}\xrightarrow{a.s.}\bm{\mu},\quad\hat{\bm{S}}_{\operatorname{MMCD}}\xrightarrow{a.s.}c(\alpha)^{-1}\bm{\Sigma}^{\operatorname{col}}\otimes\bm{\Sigma}^{\operatorname{row}}\quad,$ (45)
where $c(\alpha)>0$ is a distribution-specific consistency factor of the MCD given in Croux and Haesbroeck, (1999).
To complete the proof consider reparametrization of $l(\bm{w},\bm{a},\bm{A}|(\bm{x}_{1},\dots,\bm{x}_{n}))$ for fixed weights $\bm{w}\in\mathbb{R}^{n}$, mean $\bm{a}\in\mathbb{R}^{pq}$, and covariance $\bm{A}\in\operatorname{PDS}(pq)$ in terms of the precicion matrix $\bm{B}=\bm{A}^{-1}$. Denote this new parametrization as $g(\bm{B}|\bm{w},\bm{a},(\bm{x}_{1},\dots,\bm{x}_{n}))=l(\bm{w},\bm{a},\bm{B}^{-1}|\bm{x}_{1},\dots,\bm{x}_{n})$, which is now concave in $\bm{B}$. Especially, for $\bm{w}=\bm{w}_{\operatorname{MMCD}}$ and $\bm{a}=\hat{\bm{\mu}}_{\operatorname{MMCD}}$, the function $g(\bm{B}|\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},(\bm{x}_{1},\dots,\bm{x}_{n}))$ is concave in $\bm{B}$ and achieves a unique global maximum at $\bm{B}=\hat{\bm{S}}_{\operatorname{MMCD}}$. Equations (42) and (B.1) then give
@@
$-l(\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}}|(\bm{x}_{1},\dots,\bm{x}_{n}))<\varepsilon,$
@@
$-g((\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}})^{-1}|\bm{w}_{\operatorname{MMCD}},\hat{\bm{\mu}}_{\operatorname{MMCD}},(\bm{x}_{1},\dots,\bm{x}_{n}))<\varepsilon,$
as both $\hat{\bm{S}}_{\operatorname{MMCD}}$ and $\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}}$ are a.s. positive definite for $n$ large enough. Concavity of $g$ and the fact that $\hat{\bm{S}}_{\operatorname{MMCD}}^{-1}$ is its global maximum further imply that
$\|\hat{\bm{S}}_{\operatorname{MMCD}}^{-1}-(\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}})^{-1}\|<\delta_{1},$
for $\delta_{1}=\delta_{1}(\varepsilon)\to 0$ as $n\to\infty$. Almost sure positive definiteness of $\hat{\bm{S}}_{\operatorname{MMCD}}$ and $\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}}$, and continuity of matrix inverse imply that
$\|\hat{\bm{S}}_{\operatorname{MMCD}}-\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{col}}\otimes\hat{\bm{\Sigma}}_{\operatorname{MMCD}}^{\operatorname{row}}\|<\delta,$ (46)